OpenStudy (anonymous):

Ernesto loses 14.5 pounds in 7 months. Now he weighs 156.5 pounds. Which equation can be used to find how much he weighed before?

OpenStudy (anonymous):

A. x + 14.5 = 156.5 B. x – 14.5 = 156.5 C.7x + 14.5 = 156.5 D. 7x – 14.5 = 156.5

OpenStudy (anonymous):

B i think

OpenStudy (anonymous):

Yep :) @srw1496 is correct Weight before minus weight lost = current weight

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Shannon opens a savings account with $50. Each week after that she deposits $8 into her account. How many weeks does it take her to save $450?

OpenStudy (anonymous):

8x+50=450

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem(:

OpenStudy (anonymous):

could you tell me how many weeks it takes her to save 450

OpenStudy (anonymous):

How much extra does Shannon have to save to get from $50 to $450?

OpenStudy (anonymous):

here are the possible answers 56.25 weeks 62.5 weeks 52 weeks 50 weeks

OpenStudy (anonymous):

Well the equation I showed you is 8x+50=450. To solve for x (which is the number of weeks) you first have to get 8x by itself so you would subtract 50 from both sides and get 8x=400

OpenStudy (anonymous):

Then to get x you divide everything by 8 and get x=50. It takes her 50 weeks to save $450

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem (:

OpenStudy (anonymous):

–6q + 5.8 = 23.8

OpenStudy (anonymous):

Well done again :)

OpenStudy (anonymous):

i have a lot of math trouble if you haven't noticed

OpenStudy (anonymous):

you should post these in separate questions... then you get the "credit" for asking, plus @srw1496 can get more credit for the good help :)

OpenStudy (anonymous):

haha I have problems with math too its fine, I'll try an help as much as I can(:

OpenStudy (anonymous):

next semester i will be in an easier math class

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

–6q + 5.8 = 23.8 you need to solve for q right?

OpenStudy (anonymous):

i don't know how to get another question up other than this one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

q = 3 q = –3 q = –1.83 q = 4.93

OpenStudy (anonymous):

@cheeseloverXD it's no big deal, but if you hit the "close" button near where you asked the question, you can begin a new one.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

well you do the same thing I showed you for the last equation. You have to get the -6q on the left side alone so you have to subtract 5.8 of both sides, can you do that?

OpenStudy (anonymous):

btw thanks for the medal

OpenStudy (anonymous):

haha no problem

OpenStudy (anonymous):

how do you get a pic

OpenStudy (anonymous):

on the profile thing

OpenStudy (anonymous):

Did you work out that last problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

go to settings and click change profile picture underneath the little pic of the owl and then click save. but back to the problem now you have -6q=18

OpenStudy (anonymous):

–6q + 5.8 = 23.8 Step 1: subtract 5.8 from both sides –6q + 5.8 -5.8 = 23.8 - 5.8 Step 2: simplify: -6q = 18

OpenStudy (anonymous):

@cheeseloverXD what's the next step?

OpenStudy (anonymous):

now you have to get q by itself so you divide everything by -6 and what do you get?

OpenStudy (anonymous):

18+58?

OpenStudy (anonymous):

no, your trying to find q right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this was the only good pic I had on my comp.

OpenStudy (anonymous):

When they did the problem they multiplied -6 by q to get 18 so to get q from 18 you have to do the opposite which is to divide both -6q and 18 by -6